# black holes; early universe; general relativity

### v1

@@ 3.4 Formation threshold in radiation domination
If the amplitude of the perturbation, which is generated by inflation and enters the horizon in the radiation-dominated era, is sufficiently large, it will directly collapse to a black hole. Since the linear perturbation in the radiation-dominated era does not grow, such a perturbation must be nonlinearly large. This prevents us from directly accessing the full dynamics of PBH formation with analytical methods. Only full numerical relativity simulations can accurately describe the general relativistic dynamics of PBH formation, which has been pioneered by Nadezhin, Novikov and Polnarev (1978) Nadezhin et al. (1978). In fact, it has been established by several analytical and numerical works of the Einstein equation that PBHs really form in this scenario.
This scenario naturally implies that there exists a threshold for PBH formation. Carr (1975) derived the threshold $\delta_{th}\sim 1/3$ for radiation domination according to the Jeans scale argument in terms of $\delta_{H}$, the density perturbation at the horizon entry of the perturbation. In numerical relativity, the threshold value is obtained as $\delta_{th}\sim 0.45$ in terms of the density perturbation averaged in the comoving slice over the Hubble patch at the horizon entry of the overdense region, where $\delta_{H}$ is regarded as that in the nontrivial lowest order of the long-wavelength expansion scheme Musco et al. (2005, 2009); Musco and Miller (2013); Harada et al. (2015). Harada, Yoo and Kohri (2013) Harada et al. (2013) refined Carr’s argument from a general relativistic point of view and analytically derived the threshold $\delta_{th}\simeq 0.41$ in the comoving slice. Although the averaged density perturbation $\delta_{H}$ might look straightforward to interpret, it has difficulty in its interpretation. This is because if we calculate the averaged density perturbation at the horizon entry using the nontrivial lowest order of the solutions, it cannot be a real physical value, as the latter can only be obtained after the full numerical simulation and is not necessarily useful.
Shibata and Sasaki (1999) Shibata and Sasaki (1999) defined a compaction function in the constant-mean curvature slice. Although it was intended to equal to the ratio of the excess in the Misner-Sharp mass, or equivalently the Kodama mass, to the areal radius, it is not equal to that but to
${\cal C}_{\rm SS}:=\frac{1}{2}\left[1-\left(1+r\zeta^{\prime}\right)^{2}\right]$ (13)
in the long-wavelength limit, where we omit the subscript 00 in $\zeta$ according to the convention Harada et al. (2015, 2023). It is proportional to the ratio of the Misner-Sharp mass excess to the areal radius in the comoving slice as
@@
with $\delta M_{\rm com}$ being the Misner-Sharp mass excess in the comoving slice. Note the factor 2 on the right-hand side of Eq. (15). The threshold value is $\sim 0.4$ in terms of its maximum value of the Shibata-Sasaki compaction function ${\cal C}_{\rm SS}(r)$ in the long-wavelength limit Shibata and Sasaki (1999); Musco and Miller (2013); Harada et al. (2015). This is more straightforward than the averaged density perturbation because for the compaction function description, we only have to take the long-wavelength limit of the solutions, where all the higher-order contributions naturally disappear. This is probably why people have favoured to use the compaction function.
@@
in terms of $\bar{C}_{\rm com}$, the spatial average of $C_{\rm com}$ over $0<r<r_{m}$, where $C_{\rm com}(r)$ takes a maximum at $r=r_{m}$. This holds within $2\%$ accuracy over different profiles they surveyed. This gives a new threshold condition that uses the maximum $C_{\rm com}(r_{m})$ and the second derivative of $C^{\prime\prime}_{\rm com}(r_{m})$. See Ref. Escrivà et al. (2020) for details.
@@ 3.5 Softer equation of state
Although PBH formation has been conventionally studied in the radiation-dominated phase, the scenarios in other phases with effectively softer equations of state will also be important because the generation should be enhanced then. This scenario already was noticed in Carr (1975) Carr (1975), where the threshold density perturbation was estimated to $\delta_{H}\sim w$ for the linear EOS $p=w\rho$ using the simple application of the Jeans criterion in Newtonian gravity. Harada, Yoo and Kohri (2013) Harada et al. (2013) refined it by comparing the free-fall time and the sound-crossing time in a simplified model of the general relativistic spacetime considering gauge difference to find the threshold value
@@
in the comoving slice. This analytical expression showed a good agreement with the results of numerical relativity simulation Musco and Miller (2013); Escrivà et al. (2021) for $30<w\leq 1/$ 3 as seen in Fig. 3. These studies showed that the threshold for the EOS $p=w\rho$ is an increasing function of $w$ for $30<w\leq 1/$ 3 and approaches 0 as $w\to 0$. For example, for the QCD crossover, for which $w$ drops to $\sim 0.23$ from $1/3$ Byrnes et al. (2018), the PBH formation will be enhanced by a factor of the order of 1000 Escrivà et al. (2023); Musco et al. (2024).
Figure 3: EOS dependence of the PBH formation threshold. See also Ref. Harada et al. (2013).
@@ 3.6 Matter domination
The PBH formation in matter domination has been pioneered by Khlopov and Polnarev (1980) Khlopov and Polnarev (1980); Polnarev and Khlopov (1982), where the effects of anisotropy and inhomogeneity are studied as obstruction of PBH formation process. Harada, Yoo, Kohri, Nakao and Jhingan (2016) Harada et al. (2016) revisited the anisotropic effects by combining the picture of pancake collapse of dark matter and the hoop conjecture by Thorne, which claims that black holes with horizons form when and only when a mass $M$ gets compacted into a region whose circumference in every direction is ${\cal C}\lesssim 4\pi M$  Klauder (1972); Misner et al. (1973). This suppression due to the anisotropic effect is schematically shown in Fig. 4. They not only qualitatively reproduced the result of Ref. Khlopov and Polnarev (1980); Polnarev and Khlopov (1982) but also updated the coefficient as
$\beta_{\rm aniso}\simeq 0.05556\sigma_{H}^{5},$ (17)
where $\sigma_{H}$ is the standard deviation of $\delta_{H}$ and the Gaussian distribution for density perturbation is assumed.
Figure 4: Schematic illustration of the anisotropic effect in the formation of PBHs in matter domination. ${\cal C}$ denotes the hoop of the surface the gravitating object of mass $M$. See Ref. Harada et al. (2016) for details.
Kopp, Hofmann and Weller (2010) Kopp et al. (2011) modelled spherical formation of PBHs in matter domination with the Lemaître-Tolman-Bondi (LTB) solution and Harada and Jhingan (2015) Harada and Jhingan (2016) extended it to nonspherical formation with the Szekeres quasi-spherical solutions. Kokubu, Kyutoku, Kohri and Harada (2018) Kokubu et al. (2018) revisited the inhomogeneity effects by utilising the LTB solution and not only qualitatively reproduced the result of Ref. Khlopov and Polnarev (1980); Polnarev and Khlopov (1982) but also updated the coefficient so that the additional suppression factor is given by
$\beta_{\rm inhom}\simeq 3.70\sigma_{H}^{3/2}$ (18)
@@ 3.7 Critical behaviour
One of the great achievements of numerical relativity is the discovery of critical behaviour in gravitational collapse, which is also called black hole critical behaviour. This has been first discovered by Choptuik (1993) Choptuik (1993) in the spherically symmetric system of a massless scalar field and followed by Evans and Coleman (1994) Evans and Coleman (1994) in the spherical system of a radiation fluid. The phenomena were theoretically revealed in terms of the renormalisation group analysis by Koike, Hara and Adachi (1995) Koike et al. (1995). These early studies were all on asymptotically flat spacetimes and whether it applies also for PBH formation was not so trivial because of the different boundary conditions and the existence of the characteristic scale, the Hubble horizon length. The critical behaviour in the PBH formation was discovered by Niemeyer and Jedamzik (1999) Niemeyer and Jedamzik (1999) but subsequently questioned Hawke and Stewart (2002). Finally, Musco and Miller (2013) Musco and Miller (2013) beautifully confirmed the critical behaviour in PBH formation. The essence of the critical behaviour in the context of PBH formation is the following. Let us consider a one-parameter family of initial data, for which we can choose the averaged density perturbation $\delta_{H}$ as the parameter of the family. As we have already seen, there is a threshold value $\delta_{\rm th}$, beyond which a PBH forms. Then, the evolution of the initial data which has the critical value $\delta_{H}=\delta_{\rm th}$ approaches a particular member of continuously self-similar solutions, which is called a critical solution. If $\delta_{H}$ is slightly above the critical value, we have PBH formation and the scaling law for the mass of the formed PBH, $M$, holds as follows:
@@
where $K$ is a positive constant of the order of the unity and $\gamma\simeq 0.36$ is called a critical exponent. In fact, the critical behaviour, such as the critical solution and the critical exponent, does not depend on the choice of the one-parameter family of initial data, which is called universality. Note, however, that the critical solution and the critical exponent do depend on the matter field and the equation of state even if it is a perfect fluid Maison (1996).
In the context of PBH formation, if we assume that $\delta_{H}$ obeys some reasonable statistical distribution, the critical behaviour implies that only a tiny fraction of perturbations of mass scale $M_{H}$ can be much smaller than $M_{H}$, while a large fraction remain of the order of $M_{H}$. This becomes important especially if we consider $\beta(M)$ for $M\sim 10^{16}-10^{17}$ g. This is because the PBHs of the critical mass $\sim 10^{15}$ g are severely constrained by observation through its X-ray or gamma-ray emission, while those of $\sim 10^{16}-10^{17}$ g are not. For example, the formation probability $\beta(M)$ with the horizon mass $M=10^{16}$ g is severely constrained by such a tiny fraction of PBHs of the critical mass $\sim 10^{15}$ g considerably smaller than $10^{16}$iemeyer and Jedamzik (1998); Yokoyama (1998).
@@ 3.8 Abundance estimation and statistics
Carr (1975) Carr (1975) simply assumed that $\delta_{H}$ obeys a Gaussian distribution and obtained the following formula
@@
where $\delta_{\rm max}$ and $\sigma_{H}$ are the possible maximum value and the standard deviation of $\delta_{H}$ and in the last approximation $\delta_{\rm th}\gg\sigma_{H}$ is assumed. Since we need at least $\beta(M)\gtrsim 10^{-18}$ for PBHs to contribute to a considerable fraction of dark matter, we can conclude that only the Gaussian tail beyond $\sim 8\sigma_{H}$ is responsible for PBHs, which is shown in Fig. 5. This implies that introducing even small non-Gaussianity can significantly enhance the formation of PBHs Young and Byrnes (2013).
Figure 5: Gaussian distribution of the averaged density perturbation $\delta_{H}$. Only the tail beyond $\sim 8\sigma_{H}$ is responsible for PBHs for the typical case.
If we consider PBHs formed in the radiation-dominated era and assume $\delta_{\rm th}\simeq 0.45$, we need at least $\sigma_{H}\gtrsim 0.05$ or $P_{\zeta}(k)\gtrsim 0.01$. This is much larger than the observed value $P_{\zeta}(k)\simeq 10^{-10}$ for the CMB anisotropy, although it does not immediately exclude the considerable formation of PBHs because the scales of the PBHs and the CMB anisotropies are usually very different from each other.
Although Carr’s formula (20), which is also called the Press-Schechter approximation, is very useful, this is considered as a very rough approximation. One of the reasons is that even if the curvature perturbation obeys a Gaussian distribution, the averaged density perturbation $\delta_{H}$ in the comoving slice cannot because it must be within a finite interval between the minimum $-1$ and the maximum $2/3$ Kopp et al. (2011). The other is that since PBHs will form only at very rare peaks of the perturbation, peak theory should apply, which is known to give a physically reasonable prediction for galaxy formation Bardeen et al. (1986). The prediction of peak theory may be significantly different from that of Carr’s formula for the estimation of PBH abundance. Currently, there are a few variations in the application of peak theory to the estimation of the PBH abundance, which comes from theoretical ambiguity caused by incomplete understanding of nonlinear, nonspherical and multi-scale general relativistic dynamics Yoo et al. (2018, 2021); Germani and Sheth (2020, 2023).

### v3

@@ 3.4 Formation threshold in radiation domination
If the amplitude of the perturbation, which is generated by inflation and enters the horizon in the radiation-dominated era, is sufficiently large, it will directly collapse to a black hole. Since the linear perturbation in the radiation-dominated era does not significantly grow, such a perturbation must be nonlinearly large. This prevents us from accessing the full dynamics of PBH formation with analytical methods. Only full numerical relativity simulations can accurately describe the general relativistic dynamics of PBH formation, which has been pioneered by Nadezhin, Novikov and Polnarev (1978) Nadezhin et al. (1978). In fact, it has been established by several analytical and numerical works such as Refs. Carr (1975); Nadezhin et al. (1978) that PBHs really form in this scenario.
This scenario naturally implies that there exists a threshold for PBH formation. Carr (1975) derived the threshold $\delta_{\rm th}\sim 1/3$ for radiation domination according to the Jeans scale argument in terms of $\delta_{H}$, the density perturbation at the horizon entry of the perturbation. In numerical relativity, the threshold value is obtained as $\delta_{\rm th}\sim 0.45$ in terms of the density perturbation averaged in the comoving slice over $0<r<r_{0}$, where $r_{0}$ is the radius of the overdense region, at the horizon entry of the overdense region, where $\delta_{H}$ is regarded as that in the nontrivial lowest order of the long-wavelength expansion scheme Musco et al. (2005, 2009); Musco and Miller (2013); Harada et al. (2015). Harada, Yoo and Kohri (2013) Harada et al. (2013) refined Carr’s argument from a general relativistic point of view and analytically derived the threshold $\delta_{\rm th}\simeq 0.41$ in the comoving slice. Although the averaged density perturbation $\delta_{H}$ might look straightforward to interpret, it has difficulty in its interpretation. This is because if we calculate the averaged density perturbation at the horizon entry using the nontrivial lowest order of the solutions, it cannot be a real physical value, as the latter can only be obtained after the full numerical simulation and is not necessarily useful.
Shibata and Sasaki (1999) Shibata and Sasaki (1999) defined a compaction function in the CMC slice. Although it was intended to equal to the ratio of the excess in the Misner-Sharp mass, or equivalently the Kodama mass, to the areal radius, it is not equal to that but
${\cal C}_{\rm SS}\approx\frac{1}{2}\left[1-\left(1+r\zeta^{\prime}\right)^{2}\right],$ (15)
where the weak equality denotes the equality in the long-wavelength limit and we have omitted the subscript 00 in $\zeta$ according to the convention Harada et al. (2015, 2023). This has been shown to have a geometrical origin as a compactness function in the static spacetime obtained by removing the scale factor in the long-wavelength limit Harada et al. (2024). It is proportional to the ratio of the Misner-Sharp mass excess to the areal radius in the comoving slice as
@@
with $\delta M_{\rm com}$ being the Misner-Sharp mass excess in the comoving slice. Note the factor of 2 on the right-hand side of Eq. (17). The threshold value is $\sim 0.4$ in terms of its maximum value of the Shibata-Sasaki compaction function ${\cal C}_{\rm SS}(r)$ in the long-wavelength limit Shibata and Sasaki (1999); Musco and Miller (2013); Harada et al. (2015). This is more straightforward than the averaged density perturbation because for the compaction function description, we only have to take the long-wavelength limit of the solutions, where all the higher-order contributions naturally disappear. This is probably why people have favoured to use the compaction function. On the other hand, this threshold can be transformed to $\sim 0.5$ in terms of the density perturbation in the comoving slice averaged over the ball $0<r<r_{m}$ at its horizon entry, where $r_{m}$ is the radius at which ${\cal C}_{\rm SS}$ takes the maximum Musco (2019) again with a caveat that this value is only correct in the nontrivial lowest order of the long-wavelength expansion. The averaging over $0<r<r_{m}$ is more consistent and applicable than over the overdense ball $0<r<r_{0}$ as the definition of $r_{0}$ is problematic for density perturbations without an underdense region.
@@
in terms of $\bar{C}_{\rm com}$, the spatial average of $C_{\rm com}$ over $0<r<r_{m}$, where $C_{\rm com}(r)$ takes a maximum at $r=r_{m}$. This holds within $2\%$ accuracy over different profiles they surveyed. This gives a new threshold condition that uses the maximum $C_{\rm com}(r_{m})$ and its second-order derivative $C^{\prime\prime}_{\rm com}(r_{m})$. See Ref. Escrivà et al. (2020) for details. More recently, Ianniccari, Iovino, Kehagias, Perrone and Riotto Ianniccari et al. (2024) reported some numerical coincidence between this threshold and that for the existence of a circular photon orbit.
There is a possibility that isocurvature perturbations may be produced depending on the cosmological scenario. Yoo, Harada, Hirano, Okawa and Sasaki (2021) Yoo et al. (2022), through numerical relativity simulations by introducing a massless scalar field perturbation into the radiation-dominated Universe, demonstrated that isocurvature perturbations can also produce PBHs in the radiation-dominated era. As isocurvature perturbations can play important roles in various intriguing inflationary scenarios, further studies are anticipated to achieve a comprehensive understanding of the formation threshold from isocurvature.
@@ 3.5 Softer equation of state
There remain a lot of theoretical possibilities for the thermal history of the Universe. In addition to the standard phases of radiation dominance and subsequent matter dominance, there may be early matter-dominated phase and/or phase transition or crossover phase in which the equation of state of the dominant component of the Universe is significantly softer than the radiation fluid. Although PBH formation has been conventionally studied in the radiation-dominated phase, the scenarios in other phases with such significantly soft equations of state will also be important. This is because as we will see below, $\beta(M)$ strongly depends on the threshold, which is smaller for the softer equation of state and, hence, PBH formation can be strongly enhanced so that PBHs formed in this phase may dominate those formed in the radiation dominated era, provided that the power spectrum density perturbation is nearly scale-invariant.
This scenario was already suggested by Carr (1975) Carr (1975), where the threshold density perturbation was estimated to $\delta_{H}\sim w$ for the linear EOS $p=w\rho$ using the simple application of the Jeans criterion in Newtonian gravity. Harada, Yoo and Kohri (2013) Harada et al. (2013) refined it by comparing the free-fall time and the sound-crossing time in a simplified toy model of the fully general relativistic spacetime considering gauge difference to find the threshold value
@@
in the comoving slice. This analytical expression showed a good agreement with the results of numerical relativity simulation Musco and Miller (2013); Escrivà et al. (2021) for $30<w\leq 1/$ 3 as seen in Fig. 3. These studies showed that the threshold for the EOS $p=w\rho$ is an increasing function of $w$ for $30<w\leq 1/$ 3 and approaches 0 as $w\to 0$. For example, for the QCD crossover, for which $w$ drops to $\sim 0.23$ from $1/3$ Byrnes et al. (2018), this suggests enhanced generation of PBHs. Although $w$ is time-dependent in this case, the threshold has also been calculated based on the Jeans criterion Papanikolaou (2022). Subsequently, full numerical relativity simulations have revealed that the PBH formation will be enhanced by a factor of the order of 1000 Escrivà et al. (2023); Musco et al. (2024). See Ref. Jedamzik (2024) for a recent review on this subject.
Figure 3: EOS dependence of the PBH formation threshold. The horizontal axis is the parameter $w$ for the EOS $p=w\rho$, while the vertical axis is the threshold value in terms of $\delta_{H}$, the density perturbation over the overdense region, $0<r<r_{0}$, at its horizon entry in the comoving slice. The crosses show the numerical result obtained in Ref. Musco and Miller (2013). The purple curve is the plot of Eq. (18), which shows an agreement with the numerical result within $\sim 20\%$, while the green curve shows Carr’s formula $\delta_{H}\simeq w$. See Ref. Harada et al. (2013) for more details.
Musco and Papanikolaou (2022) Musco and Papanikolaou (2022) studied the effect of anisotropic stress, which is the deviation from the perfect fluid description, on PBH formation in radiation domination and found that if this is large enough, it could lead to a significant variation of the abundance of PBHs.
@@ 3.6 Matter domination
The PBH formation in matter domination has been pioneered by Khlopov and Polnarev (1980) Khlopov and Polnarev (1980); Polnarev and Khlopov (1981, 1982, 1985), where the effects of anisotropy and inhomogeneity are studied as obstruction of PBH formation process. Harada, Yoo, Kohri, Nakao and Jhingan (2016) Harada et al. (2016) revisited the anisotropic effects by combining the picture of pancake collapse of dark matter and the hoop conjecture by Thorne, which claims that black holes with horizons form when and only when a mass $M$ gets compactified into a region whose circumference in every direction is ${\mathscr{C}}\lesssim 4\pi M$  Klauder (1972); Misner et al. (1973). This suppression due to the anisotropic effect is schematically shown in Fig. 4. They not only qualitatively reproduced the result of Ref. Khlopov and Polnarev (1980); Polnarev and Khlopov (1982) but also updated the coefficient as
$\beta_{\rm aniso}(M)\simeq 0.05556\sigma_{H}^{5}(M),$ (19)
where $\sigma_{H}(M)$ is the standard deviation of $\delta_{H}$ in the mass scale of $M$ and the Gaussian distribution for density perturbation is assumed.
Figure 4: Schematic illustration of the scenario on the anisotropic effect in the formation of PBHs in matter domination. ${\mathscr{C}}$ denotes the hoop of the surface of the gravitating object of mass $M$. The scenario is the following. The collapse of a dust ball is unstable against nonspherical perturbations towards the pancake. Whether or not a horizon forms around the pancake is subject to the hoop conjecture. If the condition ${\mathscr{C}}\lesssim 4\pi M$ is satisfied, a black hole forms. Otherwise the collapse leads to virialisation by acquiring velocity dispersion through violent relaxation. See Ref. Harada et al. (2016) for details.
Kopp, Hofmann and Weller (2010) Kopp et al. (2011) modelled spherical formation of PBHs in matter domination with the Lemaître-Tolman-Bondi (LTB) solution and Harada and Jhingan (2015) Harada and Jhingan (2016) extended it to nonspherical formation with the Szekeres quasispherical solutions. Kokubu, Kyutoku, Kohri and Harada (2018) Kokubu et al. (2018) revisited the inhomogeneity effects by utilising the LTB solution and not only qualitatively reproduced the result of Ref. Khlopov and Polnarev (1980); Polnarev and Khlopov (1982) but also updated the coefficient so that the additional suppression factor is given by
$\beta_{\rm inhom}(M)\simeq 3.70\sigma_{H}^{3/2}(M)$ (20)
@@
There is another intriguing approach to PBH formation during the early matter-dominated era. Such an era is thought to be naturally realised during the preheating phase following inflation, when an oscillating massive scalar field dominates the Universe. It is possible to gain further insights into PBH formation during this phase by directly studying the perturbations in the Universe dominated by a massive scalar field. This approach was explicitly proposed by Padilla, Hidalgo and Malik (2022) Padilla et al. (2022, 2021); Hidalgo et al. (2023), and numerical relativity simulations have been conducted based on this method, involving a mixture of a massive scalar field and a massless scalar field de Jong et al. (2022, 2023); Aurrekoetxea et al. (2023). This approach may involve highly fruitful physics, and further studies will be necessary for a comprehensive understanding of this scenario.
@@ 3.7 Critical behaviour
One of the great achievements of numerical relativity is the discovery of critical behaviour in gravitational collapse, which is also called black hole critical behaviour. This has been first discovered by Choptuik (1993) Choptuik (1993) in the spherically symmetric system of a massless scalar field and followed by Evans and Coleman (1994) Evans and Coleman (1994) in the spherical system of a radiation fluid. The phenomena were theoretically understood in terms of the renormalisation group analysis by Koike, Hara and Adachi (1995) Koike et al. (1995). These early studies were all on asymptotically flat spacetimes and whether they apply to PBH formation was not so trivial because of the different boundary conditions and the existence of the characteristic scale, the Hubble horizon length. The critical behaviour in the PBH formation was discovered by Niemeyer and Jedamzik (1999) Niemeyer and Jedamzik (1999) but subsequently questioned Hawke and Stewart (2002). Finally, Musco and Miller (2013) Musco and Miller (2013) beautifully confirmed the critical behaviour in PBH formation. The essence of the critical behaviour in the context of PBH formation is the following. Let us consider a one-parameter family of initial data of the Cauchy problem, for which we can choose the averaged density perturbation $\delta_{H}$ as the parameter of the family. The features of the critical behaviour do not depend on the choice of the one-parameter family of the initial data. As we have already seen, there is a threshold value $\delta_{\rm th}$, beyond which a PBH forms. Then, the evolution of the initial data which has the critical value $\delta_{H}=\delta_{\rm th}$ approaches a particular member of continuously self-similar solutions, which is called a critical solution. If $\delta_{H}$ is slightly above the critical value, we have PBH formation and the scaling law for the mass of the formed PBH, $M$, holds as follows:
@@
where $\gamma\simeq 0.36$ is called a critical exponent and $K$ is a nondimensional positive constant that depends on the shape of the initial perturbation, varying between 3 and 30 with a typical value $\sim 5$. In fact, the critical behaviour, such as the critical solution and the critical exponent, does not depend on the choice of the one-parameter family of initial data, which is called universality. Note, however, that the critical solution and the critical exponent do depend on the matter field and the equation of state even if it is a perfect fluid Maison (1996).
In the context of PBH formation, if we assume that $\delta_{H}$ obeys some reasonable statistical distribution, the critical behaviour implies that a tiny fraction of perturbations of mass scale $M_{H}$ can produce PBHs that are much smaller than $M_{H}$, while a large fraction still produce those of the order of $M_{H}$. This becomes important especially if we consider $\beta(M)$ for $M\sim 10^{16}-10^{17}$ g. This is because the PBHs of the critical mass $\sim 10^{15}$ g are severely constrained by observation through its X-ray or gamma-ray emission, while those of $\sim 10^{16}-10^{17}$ g are not. For example, the formation probability $\beta(M)$ with the horizon mass $M=10^{16}$ g is severely constrained by such a tiny fraction of PBHs of the critical mass $\sim 10^{15}$ g considerably smaller than $10^{16}$iemeyer and Jedamzik (1998); Yokoyama (1998).
@@ 3.8 Abundance estimation and statistics
Carr (1975) Carr (1975); Harada et al. (2013) simply assumed that $\delta_{H}$ obeys a Gaussian distribution and obtained the following formula
@@
where $\delta_{\rm max}=2/3$ and $\sigma_{H}(M)$ are the possible maximum value and the standard deviation of $\delta_{H}$ in the mass scale of $M$ and in the last approximation $\delta_{\rm th}\gg\sigma_{H}(M)$ is assumed  [^2] The factor of two comes from the recommendation for structure formation by Press and Schechter (1975)  Press and Schechter (1974); Peebles (2020). Note that whether or not this factor applies to PBH formation is highly nontrivial.. Since we need at least $\beta(M)\gtrsim 10^{-18}$ for PBHs to contribute to a considerable fraction of dark matter, we can conclude that only the Gaussian tail beyond $\sim 8\sigma_{H}$ is responsible for PBHs. This implies that introducing small nonGaussianities can significantly enhance the formation of PBHs Young and Byrnes (2013). See Ref. Pi (2024) for a recent review on nonGaussianities in PBH formation and their link to induced gravitational waves. If we consider PBHs formed in the radiation-dominated era and assume $\delta_{\rm th}\sim 0.45$ or $\sim 0.5$, we need at least $\sigma_{H}(M)\gtrsim 0.05$ or $P_{\zeta}(k)\gtrsim 0.01$. This is much larger than the observed value $P_{\zeta}(k)\simeq 10^{-10}$ for the CMB anisotropies, although it does not immediately exclude the considerable formation of PBHs because the scales of the PBHs and the CMB anisotropies are usually very different from each other.
Although Carr’s formula (22), which is also called the Press-Schechter approximation, is very useful, this is considered as a very rough approximation. One of the reasons is that even if the curvature perturbation obeys a Gaussian distribution, the averaged density perturbation $\delta_{H}$ in the comoving slice cannot because it must be within a finite interval between the minimum $-1$ and the maximum $2/3$ Kopp et al. (2011). The other is that since PBHs will form only at very rare peaks of the perturbation, peak theory should apply, which is known to give a physically reasonable prediction for galaxy formation Bardeen et al. (1986). The prediction of peak theory may be significantly different from that of Carr’s formula for the estimation of PBH abundance. Currently, there are a few variations in the application of peak theory to the estimation of the PBH abundance, which comes from theoretical ambiguity caused by incomplete understanding of nonlinear, nonspherical and multi-scale general relativistic dynamics Yoo et al. (2018, 2021); Germani and Sheth (2020); Kitajima et al. (2021); Germani and Sheth (2023). See also Ref. Young (2024) for a recent review on this topic.